OpenStudy (anonymous):

If someone answers these for me, I will give you a medal so please, Question 1 (Multiple Choice Worth 1 points) What is the quotient: (2x2 + 12x + 18) ÷ (x + 3) ? 2x + 6 2x – 6 2x + 4 2x – 4, r = 1 Question 2 (Multiple Choice Worth 1 points) If f(x) = 3x – 1 and g(x) = x2 – 2x – 2, what is f(g(x)) ? x2 – 5x + 2 x2 + x – 3 3x2 – 6x – 7 3x2 – 5x – 3

OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 1 points) What are the possible number of positive, negative, and complex zeros of f(x) = –x6 + x5– x4 + 4x3 – 12x2 + 12 ? Positive: 4, 2, or 0; Negative: 2 or 0; Complex: 6, 4, 2, or 0 Positive: 3 or 1; Negative: 3 or 1; Complex: 4, 2, or 0 Positive: 5, 3, or 1; Negative: 1; Complex: 4, 2, or 0 Positive: 2 or 0; Negative: 2 or 0; Complex: 6, 4, or 2 Question 4 (Multiple Choice Worth 1 points) What are the possible number of positive, negative, and complex zeros of f(x) = –3x4 – 5x3 – x2 – 8x + 4 ? Positive: 2 or 0; Negative: 2 or 0; Complex: 4 or 2 or 0 Positive: 1; Negative: 3 or 1; Complex: 2 or 0 Positive: 3 or 1; Negative: 1; Complex: 2 or 0 Positive: 4 or 2 or 0; Negative: 2 or 0; Complex: 4 or 2 or 0 Question 5 (Multiple Choice Worth 1 points) According to The Fundamental Theorem of Algebra, how many zeros does the function f(x) = 17x15 + 41x12 + 13x3 – 10 have ? 15 12 30 3 Question 6 (Multiple Choice Worth 1 points) What is the quotient: (91y3 + 21y2 – 35y) ÷ 7y ? 13y2 – 3y – 3 13y2 – 3y + 5 13y3 + 3y2 – 5y 13y2 + 3y – 5 Question 7 (Multiple Choice Worth 1 points) What is the remainder when (3x4 + 2x3 – x2 + 2x – 14) ÷ (x + 2) ? 0 5 10 15 Question 8 (Multiple Choice Worth 1 points) Which of the following is a polynomial with roots –2, –2i, and 2i ? x3 – 4x2 + 4x – 16 x3 + 4x2 + 4x + 16 x3 + 2x2 + 4x + 8 x3 – 2x2 + 4x – 8 Question 9 (Multiple Choice Worth 1 points) What are the possible rational zeros of f(x) = x4 + 2x3 – 3x2 – 4x + 18 ? ± 1, ± 2, ± 3, ± 4, ± 5, ± 6, ± 7, ± 8, ± 9, ± 10, ± 11, ± 12, ± 13, ± 14, ± 15, ± 16, ± 17, ±18 ± 1, ± 2, ± 3, ± 6, ± 9, ± 18 1, 2, 3, 6, 9, 18 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18 Question 10 (Multiple Choice Worth 1 points) Which of the following is a polynomial with roots – square root of 5, square root of 5, and 3 ? x3 – 2x2 – 3x + 6 x3 + 2x2 – 3x – 6 x3 – 3x2 – 5x + 15 x3 + 3x2 – 5x – 15 Question 11 (Multiple Choice Worth 1 points) If f(x) = 5x2 – 2x + 4 and g(x) = 3x – 6, what is f(x) • g(x) ? 15x3 + 24x2 – 18x – 24 15x3 + 36x2 + 18x + 24 15x3 – 24x2 – 24x + 24 15x3 – 36x2 + 24x – 24 Question 12 (Multiple Choice Worth 1 points) What is the quotient: (6x4 – 15x3 + 10x2 – 10x + 4) ÷ (3x2 + 2) ? 2x2 – 5x + 2 2x2 + 5x – 2 2x2 – 5x – 2 2x2 + 5x + 2 Question 13 (Multiple Choice Worth 1 points) Approximate the real number solution(s) to the polynomial function f(x) = x3 – 4x2 + x + 6 . x = –3, x = –1, x = 2 x = 3, x = 1, x = –2 x = 3, x = –1, x = 2 x = –3, x = 1, x = –2 Question 14 (Multiple Choice Worth 1 points) Which of the following represent the zeros of f(x) = 6x3 + 25x2 – 24x + 5 ? –5, one–third, one–half 5, –one–third, one–half 5, one–third, – one–half 5, one–third, one–half Question 15 (Multiple Choice Worth 1 points) What are the zeros of the polynomial function: f(x) = (x + 4)(2x + 1)(x – 7)? 4, one–half, –7 –4, –one–half, 7 4, –one–half, –7 –4, one–half, 7 Question 16 (Multiple Choice Worth 1 points) If f(x) = x minus 3, all over 3, then what is f–1(x) ? one–thirdx – 1 3x – 1 3x – 3 3x + 3 Question 17 (Multiple Choice Worth 1 points) Using synthetic division, what is the quotient: (2x3 – 2x – 12) ÷ (x – 2) ? 2x2 + 4x + 6 2x2 – 4x + 6 2x2 + 4x – 6 2x2 + 4x – 6 + 1 over the quantity of x minus 2

OpenStudy (luigi0210):

you could use synthetic division for the first one

OpenStudy (anonymous):

Question 19 (Multiple Choice Worth 1 points) What is the third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3? f(x) = 4x3 – 24x2 + 22x – 24 f(x) = 2x3 – 12x2 + 22x – 12 f(x) = 4x3 + 24x2 – 44x – 24 f(x) = 2x3 + 12x2 – 22x – 12 Question 20 (Multiple Choice Worth 1 points) What are the zeros of the polynomial function: f(x) = x3 – 10x2 + 24x ? –6, 0, 4 6, 0, –4 –6, 0, –4 6, 0, 4

OpenStudy (anonymous):

2x + 6 (1st one)

OpenStudy (luigi0210):

Don't just give out answers..

OpenStudy (anonymous):

2nd one.. just plug in the g(x) function into the f(x) function \[f(g(x)) = 3(x^2-2x-2)-1\]

OpenStudy (anonymous):

its a composite function

OpenStudy (dumbcow):

holy crap ... please 1 question at a time dont just copy and paste your entire assignment, its not like you are paying for a service :P and medals do not have any value actually

OpenStudy (anonymous):

u know right.. i thought it was just 2 questions.. Now i just give up

OpenStudy (luigi0210):

Ha, thank you @dumbcow I was thinking the same thing

OpenStudy (dumbcow):

like bribing us with medals is what drives us to help people...haha